# Minimum-norm solutions of the non-symmetric semidefinite Procrustes problem

### v1

@@ 3.2 Computing and approximating the minimum-norm solution
Theorem 3.3 describes the minimum-norm and minimum-rank solution of ($\mathcal{NP}$). However an explicit solution for the problem
$\operatorname*{arg\,min}_{Y\in\mathbb{R}^{(n-r)\times s}}\|YW_{1}^{\top}-Z\|_{F}^{2}+\frac{1}{16}\|Y\Lambda^{-1}Y^{\top}\|_{F}^{2}=\operatorname*{arg\,min}_{Y\in\mathbb{R}^{(n-r)\times s}}f(Y)$ (13)
is not available.
Gradient descent
To solve (13), we can use a standard non-linear optimization algorithm, such as a gradient descent. The gradient of $f$ can be computed (see, e.g., [18]) explicitly
@@
Note that $f(Y)$ is not globally Lypschitz because of the quartic term $\|Y\Lambda^{-1}Y^{\top}\|_{F}^{2}$. However, a simple backtracking line search will work well in our case. We update the matrix $Y$ by choosing a suitable step size ${\gamma}$ such that $f(Y-{\gamma}\nabla f(Y))\leq f(Y)$, as will be discussed in detail in Algorithm 2.
Approximation via Cardano formula
In order to avoid solving the optimization problem (13), it is also possible to consider a matrix of the form
@@ 4 An algorithm for the NSPSDP problem
Algorithm 1 Algorithm for the NSPDSP problem ($\mathcal{NP}$)
@@
The first two steps of Algorithm 1 involving the solution of the reduced problem are the same as in [4]. The novelty of our algorithm is the post-processing procedure described in Steps 3, 4 and 5, which ensures that the algorithm always provides a minimum of problem ($\mathcal{NP}$), while [4] was only providing a large-norm approximate solution, see (5), in the case the infimum of ($\mathcal{NPC}$) is not attained.
@@
Algorithm 2 Backtracking line search for solving problem (13)
@@
The matrices $Z\in\mathbb{R}^{(n-r)\times r}$, $W_{1}\in\mathbb{R}^{r\times s}$, $\Lambda\in\mathbb{R}^{s\times s}$, two step size parameters $\theta_{1},\theta_{2}>1$, parameter $\underline{\gamma}$ (default=$10^{-16}$), and maximum number of iterations maxiter (default=200).
@@
1:Initialize $Y_{0}=\alpha_{\diamond}ZW_{1}$ defined by the Cardano approximation (14) and set
$\gamma_{0}=\frac{\|Y_{0}\|_{F}^{2}}{\|\nabla f(Y_{0})\|_{F}^{2}}.$
2:for $i=0,1,\dots,$ maxiter do
3:     Compute
$\nabla f(Y_{i})=\frac{1}{4}Y_{i}\Lambda^{-1}Y_{i}^{\top}Y_{i}\Lambda^{-1}+2Y_{i}-2ZW_{1}.$
4:     Set $\gamma_{i}^{(0)}=\gamma_{i}$ and $k=0$.
5:     while $f(Y_{i}-\gamma_{i}\nabla f(Y_{i}))>f(Y_{i})$ and $\gamma_{i}^{(k)}>\underline{\gamma}$ do
6:         $\gamma_{i}^{(k+1)}=\gamma_{i}^{(k)}/\theta_{1}$.
7:         $k:=k+1$.
8:     end while
9:     Set $Y_{i+1}=Y_{i}-\gamma_{i}^{(k)}\nabla f(Y_{i})$ and update $\gamma_{i+1}:=\gamma_{i}^{(k)}\theta_{2}$.
10:end for
The parameters $\theta_{1}$ and $\theta_{2}$ introduced in Algorithm 2 determine the enlargement and shrinking rate for the step size $\gamma$. In our experiments we have always used $\theta_{1}=2$ and $\theta_{2}=1.1$. In Section 5 we give more details about the stopping criterion for the algorithm.
Approximate solution with Cardano
@@
has a symmetric part of rank $s$ and skew-symmetric part of rank bounded by $2r$, as shown in the proof of Theorem 3.3 for $Y_{\star}$. However the solution $A_{\diamond}$ is not guaranteed to have minimum Frobenius norm. In the numerical experiments in Section 5, we will compare these two approaches, both in term of computational time and size of the norm of the solution.
@@ 4.1 Computational complexity
    As shown in [4], the complexity of the fast gradient method applied to the $r\times r$ NSPDSP problem is dominated by the computation of the gradient $AXX^{\top}-BX^{\top}$, which is $\mathcal{O}(\tau_{1}r^{3})$, where $\tau_{1}$ is the number of iterations performed. See Section 5 below for a discussion on the stopping criterion used.
@@ 4.2 The algorithm for nearly low-rank $X$
$\tilde{\gamma}:=\min_{A\in\mathcal{N}^{n}_{\succeq}}\|A\tilde{X}-B\|_{F}=\|\tilde{A}\tilde{X}-B\|_{F},\qquad\gamma_{\star}:=\min_{A\in\mathcal{N}^{n}_{\succeq}}\|AX-B\|_{F}=\|A_{\star}X-B\|_{F},$
where $A_{\star}$ is the unique minimizer of the original NSPDSP problem, while $\tilde{A}$ is a minimizer of the perturbed NSPDSP problem with smallest norm. Then
@@ 5 Numerical experiments
Regarding Algorithm 2 for the computation of $Y_{\star}$ in MINGD, we stop the gradient descent algorithm when the iterate $Y^{(k)}$ satisfies the relative error bound
$\frac{\|Y^{(k)}-Y^{(k-1)}\|_{F}}{\|Y^{(k)}\|_{F}}\leq 10^{-8},$
or when the number of iterations exceeds 200.
@@ 5.1 Synthetic data
MINGD $0.8674\pm 0.0114$ $\mathbf{0.9331\pm 0.0575}$ $\mathbf{<0.01}$ $\mathbf{9.40\pm 0.68}$ $40\pm 0$
@@
MINGD $0.8007\pm 0.0060$ $\mathbf{1.1859\pm 0.0435}$ $\mathbf{0.03\pm 0.01}$ $\mathbf{18.90\pm 0.64}$ $80\pm 0$
@@
MINGD $0.8776\pm 0.0028$ $\mathbf{0.8145\pm 0.0225}$ $\mathbf{0.03\pm 0.00}$ $\mathbf{23.95\pm 0.83}$ $100\pm 0$
@@
  • All the methods provide the same relative error, up to a difference of order $10^{-9}$. Although ANFGM does not solve the same problem, it constructs a solution using the parameter $\varepsilon=10^{-8}$ in (5), and hence the solution generated has a relative error very close to that of the other algorithms.
@@
In summary the best methods are MINGD and CARD: they provide a solution with small norm (the one from CARD is generally $2\%$ larger than that of MINGD), the computation is fast (MINGD is a bit slower than CARD; see the next sections for more experiments) and they satisfy the low-rank properties of Theorem 3.3, which makes the storage of the solution more efficient. Instead FGM is quite slow, while ANFGM provides solutions with very large norm and neither of these two methods provides a solution whose symmetric part has the smallest possible rank.
@@ 5.3 A large low-rank example
and we solve the NSPDSP problem ($\mathcal{NP}$) for $X=J_{n}J_{m}^{\top}$ and $B=H$. The results are shown in Table 5.7.
@@
Table 5.7: Results for the matrix NSPDSP problem with $X=J_{n}J_{m}^{\top}$ and $B=H$. Best solution highlighted in bold, if applicable.
We observe similar results as the ones obtained for the synthetic data. In particular all the methods have the same relative error, ANFGM provides a large norm solution, and FGM is by far the slowest method. In contrast MINGD and CARD perform well and their solutions have low-rank properties. This example shows that it is possible to solve the low-rank NSPDSP problem of large dimension relatively fast, we further discuss the scalability in the next section.
@@ 5.4 Scalability
Figure 5.1 shows a similar trend for both examples. For ANFGM, MINGD and CARD, the complexity is dominated by the SVD computation ($\mathcal{O}(mn^{2})$ flops) and we notice that the trend is approximately a power function of $n$ with exponent between $2$ and $3$. Instead, the FGM is significantly more expensive and its time requirement is larger for $n=2^{9}$ than that of the other three method performed on dimension $n=2^{12}$.
We also observe that in the second example the MINGD and CARD perform better than ANFGM for large values of $n$ and in particular CARD turns out to be the fastest algorithm.
@@ 6 Conclusion
In this paper, we proposed a state-of-the-art semi-analytical approach for the NSPDSP problem. Our approach is inspired by that of Baghel et al. [4], but we resolved an issue when $X$ is rank deficient. By doing so, we were able to prove that the solution of the NSPDSP problem is always attained (Theorem 3.1). Moreover, we proposed a way to compute the minimum-norm and minimum-rank solution (Theorem 3.3). We illustrated the effectiveness of the new proposed algorithm on several numerical experiments.

### v2

@@ 3.2 Computing and approximating the minimum-norm solution
Theorem 3.4 describes the minimum-norm and minimum-rank solution of ($\mathcal{NP}$). However a closed-form for the unique solution (Lemma 3.3) of the problem,
$Y_{\star}=\operatorname*{arg\,min}_{Y\in\mathbb{R}^{(n-r)\times s}}\|YW_{1}^{\top}-Z\|_{F}^{2}+\frac{1}{16}\|Y\Lambda^{-1}Y^{\top}\|_{F}^{2}=\operatorname*{arg\,min}_{Y\in\mathbb{R}^{(n-r)\times s}}f(Y),$ (13)
is not available. We propose two different approaches to overcome this issue.
Computing the minimum-norm solution
To solve (13), we use a standard non-linear optimization algorithm, namely gradient descent. The gradient of $f$ can be computed (see, e.g., [21]) explicitly:
@@
Note that $f(Y)$ is not globally Lipschitz because of the quartic term $\|Y\Lambda^{-1}Y^{\top}\|_{F}^{2}$. Although gradient descent is a relatively naive method, it worked very well for our purpose, because the problem is typically well conditioned (the Hessian is the identity plus a semidefinite matrix). In particular a simple backtracking line search will work well in our case. We update the current approximation $Y$ by choosing a suitable step size ${\gamma}$ that satisfies the Armijo condition:
$f(Y-\gamma\nabla f(Y))\leq f(Y)-c\gamma\|\nabla f(Y)\|^{2},$ (14)
where $10<c<$ 1. This allows us to get global convergence towards $Y_{\star}$, since $f$ is trivially bounded from below by 00 and $\nabla f$ is locally Lipschitz continuous (see [20, Theorem 3.2]). We will provide the details in Algorithm 2.
Replacing the minimum-norm solution using Cardano’s formula
Instead of approximating $Y_{\star}$ using gradient descent applied on (13), we propose the following approximation:
@@ 4 An algorithm for the NSPSDP problem
Algorithm 1 Algorithm for the NSPSDP problem ($\mathcal{NP}$)
@@
The first two steps of Algorithm 1 involving the solution of the reduced problem are the same as in [4]. The novelty of our algorithm is the post-processing procedure described in Steps 3, 4 and 5, which ensures that the algorithm always provides a minimum of problem ($\mathcal{NP}$), while [4] was only providing a large-norm approximate solution (see (5)), in the case the infimum of ($\mathcal{NPC}$) is not attained.
@@
Algorithm 2 Backtracking line search with Armijo condition for solving problem (13)
@@
The matrices $Z\in\mathbb{R}^{(n-r)\times r}$, $W_{1}\in\mathbb{R}^{r\times s}$, $\Lambda\in\mathbb{R}^{s\times s}$, a stepsize parameter $\theta>1$, a parameter $c>0$, maximum number of iterations $i_{\max}$, maximum number of checks for the Armijo condition $j_{\max}$ and an error tolerance $\tau$
@@
1:Initialize $Y_{0}=\alpha_{\diamond}ZW_{1}$ defined by the Cardano approximation (15).
2:Set $i=1$ and $e_{0}=1$; compute $f(Y_{0})$ and
$\nabla f(Y_{0})=\frac{1}{4}Y_{0}\Lambda^{-1}Y_{0}^{\top}Y_{0}\Lambda^{-1}+2Y_{0}-2ZW_{1}.$
3:while $i<i_{\max}$ and $\|\nabla f(Y_{i})\|_{F}\leq\tau$ and $e_{i}\leq 10^{-15}$ do
4:     Set
$\gamma=\frac{1}{10}\cdot\frac{\|Y_{i}\|_{F}}{\|\nabla f(Y_{i})\|_{F}}.$
5:     for $j=1,\dots,j_{\max}$ do
6:         Compute $\widehat{Y}=Y_{i}-\gamma\nabla f(Y_{i})$.
7:         if $f(\widehat{Y})\leq f(Y_{i})-c\gamma\|\nabla f(Y_{i})\|_{F}^{2}$ then
8:              break
9:         else
10:              $\gamma=\gamma/\theta$.
11:         end if
12:     end for
13:     Update $Y_{i+1}:=\hat{Y}$ and compute $\nabla f(Y_{i+1})$.
14:     Set $i=i+1$ and $e_{i}=\frac{|f(Y_{i+1})-f(Y_{i})|}{f(Y_{i})}$.
15:end while
16:return $Y_{\star}:=Y_{i}$.
The parameter $\theta$ introduced in the algorithm determines the reduction rate for the stepsize $\gamma$, the constant $c$ is used in the Armijo condition (14), and the safety parameter $j_{\max}=50$ stops the algorithm in the case the line search fails to fulfill (14). In our experiments, we have always used $\theta=1.5$ and $c=10^{-4}$. In Section 5, we give more details about the stopping criterion for the algorithm.
Computing $Y_{\diamond}$
@@
has a symmetric part of rank $s$ and skew-symmetric part of rank bounded by $2r$, as shown in the proof of Theorem 3.4 for $Y_{\star}$. However the solution $A_{\diamond}$ is not guaranteed to have minimum Frobenius norm. In the numerical experiments in Section 5, we will compare these two approaches, both in terms of computational time and size of the norm of the solution.
@@ 4.1 Computational complexity
    As shown in [4], the complexity of the fast gradient method applied to the $r\times r$ NSPSDP problem is dominated by the computation of the gradient $AXX^{\top}-BX^{\top}$, which is $\mathcal{O}(\tau_{1}r^{3})$, where $\tau_{1}$ is the number of iterations performed. See Section 5 below for a discussion on the stopping criterion used.
@@ 4.2 The algorithm for nearly low-rank $X$
$\tilde{\gamma}:=\min_{A\in\mathcal{N}^{n}_{\succeq}}\|A\tilde{X}-B\|_{F}:=\|\tilde{A}\tilde{X}-B\|_{F},\qquad\gamma_{\star}:=\min_{A\in\mathcal{N}^{n}_{\succeq}}\|AX-B\|_{F}:=\|A_{\star}X-B\|_{F},$
where $A_{\star}$ is the unique minimizer of the original NSPSDP problem, while $\tilde{A}$ is a minimizer of the perturbed NSPSDP problem with smallest norm. Then
@@ 5 Numerical experiments
Regarding Algorithm 2 for the computation of $Y_{\star}$ in MINGD, we stop the gradient descent algorithm when the iterate $Y^{(k)}$ satisfies one of the following conditions
$\|\nabla(Y^{(k)})\|_{F}\leq\tau:=10^{-8},\qquad\frac{|f(Y^{(k+1)})-f(Y^{(k)})|}{f(Y^{(k)})}\leq 10^{-15},$
or when the number of iteration number exceeds $i_{\max}=500$. The stopping criterion that is most often reached first is the one based on the norm of the gradient. However, in some cases, we observed that the objective function reached machine accuracy (about $10^{-15}$) while the norm of the gradient was significantly larger than $10^{-8}$. The reason is the presence of a quartic term in $f$.
@@ 5.1 Synthetic data
MINGD $0.8674\pm 0.0114$ $\mathbf{0.9116\pm 0.0576}$ $\mathbf{<0.01}$ $\mathbf{9.40\pm 0.68}$ $40\pm 0$
@@
MINGD $0.8007\pm 0.0060$ $\mathbf{1.1459\pm 0.0285}$ $0.05\pm 0.03$ $\mathbf{18.90\pm 0.64}$ $80\pm 0$
@@
MINGD $0.8776\pm 0.0028$ $\mathbf{0.7859\pm 0.0154}$ $0.12\pm 0.08$ $\mathbf{23.95\pm 0.83}$ $100\pm 0$
@@
  • All the methods provide the same relative error, up to a difference of order $10^{-9}$. Although ANFGM does not solve the same problem, it constructs a solution using the parameter $\varepsilon=10^{-8}$ in (5), and hence the solution generated has a relative error very close to that of the other algorithms, even though MINGD and CARD are preferable since they attain a minimum.
@@
In summary the best methods on our set of synthetic data are MINGD and CARD: they provide a solution with small norm (the one from CARD is generally $4\%$ larger than that of MINGD), the computation is fast (MINGD is a bit slower than CARD; see the next sections for more experiments) and they satisfy the low-rank properties of Theorem 3.4, which makes the storage of the solution more efficient. Instead FGM is quite slow, while ANFGM provides solutions with very large norm and neither of these two methods provides a solution whose symmetric part has the smallest possible rank.
@@ 5.3 A large low-rank example
and we solve the NSPSDP problem ($\mathcal{NP}$) for $X=J_{n}J_{m}^{\top}$ and $B=H$. The results are shown in Table 5.7.
@@
Table 5.7: Results for the matrix NSPSDP problem with $X=J_{n}J_{m}^{\top}$ and $B=H$. Best solution highlighted in bold, if applicable.
We observe similar results as the ones obtained for the synthetic data. In particular all the methods have the same relative error, ANFGM provides a large norm solution, and FGM is by far the slowest method. In contrast MINGD and CARD perform well and their solutions have low-rank properties. This example shows that it is possible to solve the low-rank NSPSDP problem of large dimension relatively fast, we further discuss the scalability in the next section.
5.4 A real-world application
As a practical example, we study the local compliance estimation problem already considered in [14, 4], concerning a tiger plush toy. The dimensions of the problem are $n=3$ and $m=12$ and the data matrices (with two digits of accuracy) are
$X^{\top}=\begin{pmatrix}-0.32&0.03&0.06\\ -0.33&-0.02&0.06\\ -0.36&0.08&0.06\\ -0.30&0.03&0.05\\ -0.32&0.00&0.07\\ -0.34&0.07&0.05\\ -0.24&0.07&0.05\\ -0.21&-0.01&0.02\\ -0.33&0.16&0.10\\ -0.25&0.09&0.06\\ -0.22&0.00&0.03\\ -0.31&0.15&0.09\end{pmatrix},\qquad B^{\top}=\begin{pmatrix}-1.43&0.15&-0.44\\ -1.40&-0.31&-0.42\\ -1.38&0.44&-0.42\\ -1.43&0.14&-0.44\\ -1.40&-0.31&-0.42\\ -1.37&0.43&-0.42\\ -1.43&0.16&-0.43\\ -1.40&-0.32&-0.42\\ -1.38&0.42&-0.43\\ -1.43&0.15&-0.44\\ -1.40&-0.33&-0.42\\ -1.37&0.42&-0.44\end{pmatrix}.$
The matrix $X$ has rank $r=n=3$, hence the solution to the problem is unique and all the methods converge to the same matrix provided by [4, 14], with a relative error of $18.99\%$. In order to run the algorithms in a non-full rank setting, we replace $X$ by its best rank-2 approximation, $X_{2}$. Note that $\|X-X_{2}\|_{F}/\|X\|_{F}=3.5\%$ so $X_{2}$ is a good approximation of $X$. Table 5.8 shows the results of the methods applied on $X_{2}$.
ANFGM (lr) 0.2022 6.6785 <0.01 2 2
FGM (lr) 0.2022 7.4655 0.0158 3 2
MINGD (lr) 0.2022 6.6781 0.0143 2 2
CARD (lr) 0.2022 6.6781 <0.01 2 2
Table 5.8: Results for the tiger plush toy example where $X$ is replaced by its best rank-$2$ approximation, $X_{2}$.
All methods find a solution with minimum relative error. However, CARD and MINGD are the only ones to find a minimizer with minimum norm and minimum rank; FGM does not find a solution with minimum rank nor minimum norm, and ANFGM does not find a solution with minimum norm.
@@ 5.5 Scalability
Figure 5.1 shows a similar trend for both examples. For ANFGM, MINGD and CARD, the complexity is dominated by the SVD computation ($\mathcal{O}(mn^{2})$ flops). The trend of the computational time is approximately a power function of $n$ with exponent between $2$ and $3$. Also in these cases ANFGM is generating solutions with very large norms (e.g., larger than $10^{14}$ for $j\geq 9$), as in the previous examples, while FGM is significantly more expensive and its time requirement is larger for $n=2^{9}$ than that of the other three methods in dimension $n=2^{11}$.
Interestingly, for the second example (Figure 5.1, right plot), MINGD and CARD are faster than ANFGM for large values of $n$ (namely $n=2^{11},2^{12}$). The reason is the initialization step of ANFGM that does not scale as well. For $n=2^{12}$, the solution of MINGD and CARD coincide, and their CPU time is the same.
@@ 6 Conclusion
In this paper, we proposed a state-of-the-art semi-analytical approach for the NSPSDP problem. Our approach is inspired by that of Baghel et al. [4], but we resolved an issue when $X$ is rank deficient. By doing so, we were able to prove that the solution of the NSPSDP problem is always attained (Theorem 3.1). Moreover, we proposed a way to compute the minimum-norm and minimum-rank solution (Theorem 3.4). We illustrated the effectiveness of the new proposed algorithm on several numerical experiments.
Acknowledgment
We are very grateful to the editor and the anonymous reviewers who carefully read the manuscript; their feedback allowed us to improve it significantly.